OpenStudy (anonymous):

Hello i need help plz What is the LCD of 1/3 and 3/4 ? A. 3 B. 4 C. 7 D. 12

OpenStudy (anonymous):

What is the LCD of 1/3 and 3/4 ? A. 3 B. 4 C. 7 D. 12

OpenStudy (anonymous):

what are the denominators? are they relatively prime?

OpenStudy (anonymous):

@YoungStudier

OpenStudy (anonymous):

just looking for answers?

OpenStudy (anonymous):

idk sorry

OpenStudy (anonymous):

no

OpenStudy (anonymous):

looking to see what the answer is and why thats the answer

OpenStudy (anonymous):

i dont allow direct answer

OpenStudy (anonymous):

is the answer B

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

least common denominator is 3*4=?

OpenStudy (anonymous):

i12

OpenStudy (alexandervonhumboldt2):

4 is not made up of 3 and something. yes it is 12

OpenStudy (anonymous):

so d?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

What is the LCD of 3/10 and -4/15 ?

OpenStudy (anonymous):

is it 3*15?

OpenStudy (alexandervonhumboldt2):

no it is not 3*15 look: 10=5*2 15=5*3 => least common denominator is 5*2*3=?

OpenStudy (alexandervonhumboldt2):

you need to write a prime factorization of denominators

OpenStudy (anonymous):

so 30?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

negative or a positive?

OpenStudy (anonymous):

file:///Users/Thaqib/Documents/K12%20sheet.html

OpenStudy (anonymous):

copy and paste on to another tab

OpenStudy (anonymous):

@Xyeana